OpenStudy (anonymous):

6x-2y=32 4x-2y=18 solve?

OpenStudy (anonymous):

can you make it so that one of these equations just looks like this? y = ____

OpenStudy (mathmate):

Subtract the second equation from the first and solve for x

OpenStudy (anonymous):

mathmate's approach would work as well you can subtract them and then you will know what x will be

OpenStudy (anonymous):

i'm sorry but can someone walk me through?

OpenStudy (mathmate):

ok, let's start by subtracting the second equation from the first: 6x-2y=32 ...(1) 4x-2y=18 ...(2) (1)-(2) 6x-4x=32-18, => 2x=14 => x=7 Now put x=7 in equation (1) 6(7)-2y=32 42-2y=32 2y=42-32=10 y=10/2=5 So x=7, y=5

OpenStudy (anonymous):

hm

OpenStudy (mathmate):

Is everything ok?